OpenStudy (billyjean):

determine whether 8x+1 and 4x+1 are non inverse or inverse?

OpenStudy (3mar):

May I help?

OpenStudy (billyjean):

yes

OpenStudy (3mar):

With my pleasure! which one do you want to start to get the other one?

OpenStudy (billyjean):

it doesnt matter

OpenStudy (3mar):

You choose one!

OpenStudy (billyjean):

f(X)=5x+1/x

OpenStudy (3mar):

Great! Do you know how to get the inverse of a function?

OpenStudy (billyjean):

no

OpenStudy (billyjean):

im kind of confused

OpenStudy (3mar):

No problem at all! Step by step, it will be easy In Sha' Allah To get the inverse of any function f(x) [ just make sure that it passes the horizontal line test] follow the steps: - Solve for x, i.e separate x in one side and the other terms in the other side, included f(x). - Swap x and f(x). - The result is the inverse of the original function \(f^{-1}(x)\). try these!

OpenStudy (billyjean):

im still kind of confused of how to find the f(x)=5x+1/x for inverse functions

OpenStudy (3mar):

Let's do it together! agree?

OpenStudy (billyjean):

yup

OpenStudy (3mar):

Great! \[\LARGE f(x)=\frac{ 5x+1 }{ x }\] Cross multiply: \[\LARGE x*f(x)=5x+1\] Make all x-terms in one side: \[\LARGE x*f(x)-5x=1\] Take x as a common factor: \[\LARGE x[f(x)-5]=1\] Solve for x now: \[\LARGE x=\frac{ 1 }{ f(x)-5 }\] Swap x and f(x): \[\LARGE f^{-1}=\frac{ 1 }{ x-5 }\] This is the inverse of f(x)!

OpenStudy (3mar):

@billyjean Do you follow?

OpenStudy (billyjean):

yes

OpenStudy (billyjean):

what about for g(x)=x/5x+1 im kind of confused for that one

OpenStudy (3mar):

The same idea of the same procedure0! Can you do it!

OpenStudy (billyjean):

sorry im really confused on this question

OpenStudy (billyjean):

i really need some help

OpenStudy (3mar):

So Where are you stuck?

OpenStudy (billyjean):

for the inverse of x/5x+1 I got x/5x-1

OpenStudy (3mar):

Not quite correct! Let's represent it as follows: \[\LARGE\color{SeaGreen} {g(x)=\frac{x}{5x+1}}\] Cross multiply: \[\LARGE\color{SeaGreen} {g(x)*[5x+1]=x}\] Distribute g(x): \[\LARGE\color{SeaGreen} {5x*g(x)+g(x)=x}\] Gather all x-terms in one side: \[\LARGE\color{SeaGreen} {5x*g(x)-x=-g(x)}\] Take x as a common factor: \[\LARGE\color{SeaGreen} {x[5*g(x)-1]=-g(x)}\] Solve for x: \[\LARGE\color{SeaGreen} {x=\frac{ -g(x) }{ 5*g(x)-1 }}\] or \[\LARGE\color{SeaGreen} {x=\frac{ g(x) }{ 1-5*g(x)}}\] Swap x and g(x): \[\LARGE\color{DarkGreen } {g^{-1}(x)=\frac{ x }{ 1-5x}}\] \[\large\underline{\color{MediumBlue }{ Are ~you~ persuaded~?}}\]

OpenStudy (3mar):

@billyjean Do you follow?

OpenStudy (billyjean):

yes

OpenStudy (3mar):

I am happy to hear that.

OpenStudy (3mar):

any difficulties?

OpenStudy (billyjean):

no thank you so much

OpenStudy (3mar):

Don't mention it! Any Help... Any Time... That is with my pleasure!

OpenStudy (3mar):

Any more questions?

OpenStudy (billyjean):

no

OpenStudy (3mar):

I will not be late for any help.

OpenStudy (solomonzelman):

In general \(\color{black}{\displaystyle f }\) and \(\color{black}{\displaystyle g }\) are inverses of each other, iff \(\color{black}{\displaystyle f[g(x)]=g[f(x)]=x }\).

OpenStudy (solomonzelman):

Alternatively, you can try to find \(\color{black}{\displaystyle f^{-1}(x)}\) or \(\color{black}{\displaystyle g^{-1}(x)}\) , and if it turns out that \(\color{black}{\displaystyle f^{-1}(x)\ne g(x)}\) or \(\color{black}{\displaystyle g^{-1}(x)\ne f(x)}\), then, they are not inverses.

OpenStudy (solomonzelman):

So, by this second approach, all you have to do is to find the inverse function of either \(f\) or \(g\).